OpenStudy (anonymous):

The graph of a limacon curve is given. Without using your graphing calculator, determine which equation is correct for the graph. [-5, 5] by [-5, 5]

OpenStudy (anonymous):



OpenStudy (anonymous):

the answer choices are; r = 2 - 2 cos θ r = 3 - 2 cos θ r = 3 - cos θ r = 3 - sin θ